# Bag of Design Choices for Inference of High-Resolution Masked Generative Transformer

### v1

@@ Abstract
Text-to-image diffusion models (DMs) develop at an unprecedented pace, supported by thorough theoretical exploration and empirical analysis. Unfortunately, the discrepancy between DMs and autoregressive models (ARMs) complicates the path toward achieving the goal of unified vision and language generation. Recently, the masked generative Transformer (MGT) serves as a promising intermediary between DM and ARM by predicting randomly masked image tokens (i.e., masked image modeling), combining the efficiency of DM with the discrete token nature of ARM. However, we find that the comprehensive analyses regarding the inference for MGT are virtually non-existent, and thus we aim to present positive design choices to fill this gap. We modify and re-design a set of DM-based inference techniques for MGT and further elucidate their performance on MGT. We also discuss the approach to correcting token’s distribution to enhance inference. Extensive experiments and empirical analyses lead to concrete and effective design choices, and these design choices can be merged to achieve further performance gains. For instance, in terms of enhanced inference, we achieve winning rates of approximately 70% compared to vanilla sampling on HPS v2 with the recent SOTA MGT Meissonic. Our contributions have the potential to further enhance the capabilities and future development of MGTs.
@@ 1 Introduction
Figure 2: The complete sampling pipeline of MaskGIT, Muse, and Meissonic, and how our proposed effective and specific design choices integrate into that sampling pipeline. Specifically, TomeMGT and SCQ act on the multi-model (MM) Transformer to reduce inference latency and memory usage, respectively. Meanwhile, noise regularization and differential sampling enhance inference by correcting the probability distribution applied for sampling. Additionally, masked Z-Sampling is a rescheduling technique that significantly improves the quality of synthesized images through the forward-inversion operator (i.e., sampling and backtracking alternatively).
@@
To fill this gap, this paper focuses on the inference phase of MGT, aiming to identify effective design choices for enhancing visual quality and achieving efficient sampling using empirical analysis in high-resolution image synthesis scenarios. First, we elucidate how well-known training-free methods in DM can be applied to MGT, and the redesign required to ensure their effectiveness. As illustrated in Fig. 2 and Table 1, the sampling process of MGT bears significant similarity to that of DM, making it intuitive to adapt those algorithms from DM to MGT. We explore DPM-Solver [24, 23], TomeSD [4], and Z-Sampling [2] in this context, but find that all three algorithms require specific modifications to align with the characteristics of MGT in order to reduce NFE, accelerate inference, or achieve performance improvements. Given an example of Z-Sampling, we find that implementing DDIM Inversion [26], as used in DM, with random masking is ineffective unless masking is limited to low-confidence predicted tokens. In particular, our experimental outcomes indicate that among these three algorithms, DPM-Solver and TomeSD have relatively limited effects on MGT, whereas a rescheduling algorithm like Z-Sampling yields notable performance gains. Furthermore, we investigate the noise schedule in MGT, akin in EDM [14], highlighting that the cosine noise schedule is suboptimal under certain conditions. These findings suggest that inconsistencies between the training and inference mechanisms of DM and MGT lead to enhanced inference algorithms on DM not necessarily being effective for MGT.
@@
Second, we take a look at the probability distribution generated by the backbone (i.e., Transformer), leading to the development of several “cheap” (i.e., w/o significant computational overhead) yet effective distribution correction algorithms, including noise regularization and (low-entropy) differential sampling. To be specific, noise regularization dynamically applies (Gaussian) perturbations to the backbone output based on timesteps before applying softmax, aiming to enhance the diversity of synthesized images. Differential sampling, on the other hand, calculates the Kullback-Leibler (KL) divergence between the outputs of two adjacent time steps and resamples tokens with excessively similar Transformer outputs, thereby avoiding information redundancy and enhancing the visual quality.
@@
Noise Schedule (Section 3.1) $\sigma_{t}$ $\sqrt{e^{a^{2}t+bt}-1}$ (VP-SDE [39]) $t$ (flow matching [22]) $(\sigma_{\textrm{max}}^{\frac{1}{\rho}}+t(\sigma_{\textrm{min}}^{\frac{1}{\rho}}-\sigma_{\textrm{max}}^{\frac{1}{\rho}}))^{\rho}$ (EDM [14]) N/A, and predicts one token per iteration $\mathrm{cos}\left(\frac{\pi t}{2}\right)$ $(1-t)^{\rho}$ or $1-t^{\rho}$
@@
Fourth, we validate the effectiveness of our proposed design choices on the HPD v2 benchmark using various metrics, including ImageReward [50], HPS v2 [48], PickScore [10, 19], and AES [17]. As illustrated in Fig. 1, these strategies significantly improve the visual quality of synthesized images. Furthermore, we introduce Challengebench, which includes 220 prompts to evaluate the robustness of Meissonic [3] on challenging cases. The results demonstrate that Meissonic shows a greater performance gain on Challengebench compared to SD XL [29], as opposed to the HPD v2 benchmark.
@@ 3.3 Noise Regularization
where $\mathbf{v}_{i}$ is the output of $f_{\theta}$ and the yellow part stands for noise regularization. As shown in Fig. 7 (Left), vanilla sampling often results in horizontal bars of the same color as $t$ approaches 0 and vertical bars of the same color as $t$ approaches 1, indicating the redundancy of its sampling process. Noise regularization introduces randomness into the sampling process, eliminating color bars of the same color and thereby enhancing the diversity of predicted tokens.
@@
In our definition, the hyperparameter $\mathbf{I}_{t}$ is a function $\mathbf{I}(t)$ that varies dynamically with $t$. To empirically determine the empirically optimal $\mathbf{I}(t)$, we test various curves and calculate the winnowing rates of four metrics—PickScore, HPS v2, AES, and ImageReward—relative to vanilla sampling. The results are presented in Fig. 8. We find that $\mathbf{I}(t)$ works best when set to $|\cos(\pi t)|$, which reaches larger values as $t$ approaches 00 or $1$ and lower values in the middle. This observation is interesting, as injecting noise during the end-of-sampling phase significantly degrades inference performance on DM, while having positive impact on MGT. Additionally, as presented in Fig. 7 (Right), noise regularization modifies the entropy of the distribution during the sampling phase, making it more inclined toward a “U”-shaped structure, which is similar to the shape of $|\cos(\pi t)|$. The results on HPD v2 can be found in Appendix 9.3.
@@ 3.4 Differential Sampling
Model Method HPS v2 (↑) ImageReward (↑) AES (↑) PickScore (↑)
Meissonic N/A 0.2116 0.0670 5.8503 19.3237
SD XL N/A 0.1838 -0.4915 5.8195 19.5792
FLUX.1-schnell (quant 8bit) N/A 0.2257 0.2595 6.0357 19.6465
SD-3.5-Large (quant 4bit) N/A 0.2114 0.0373 6.0848 19.7922
Table 4: Comparison of various models on Challengebench. The result of our methods can be found in Appendix 9.8.
@@ 4.2 Introducing TomeSD into MGT
Applying token merging to accelerate inference is natural, as the backbone of MGT is a transformer. Unfortunately, Meissonic has only 1024 tokens, fewer than 4096 tokens in the attention layer of SD XL. As is well known, the computational complexity of attention increases exponentially with the number of tokens, and a smaller number of tokens reduces the potential benefits of token merging, resulting in an insignificant effect of our implemented TomeMGT as observed in our experiments. Therefore, we focus on the challenges of applying TomeSD, which has demonstrated effective on SD XL, to MGT to achieve accelerated inference, along with the corresponding application scenarios.
@@ 4.3 Momentum-based Solver
5 Challengebench and Additional Evaluations
We further analyze the performance of MGT on challenging prompts. We synthesize 150k images using SD XL [29], computed HPS v2 scores, and selected 1.5k prompts with the lowest scores. After manual and GPT-4o [1] selection, we form Challengebench with 220 semantically sound prompts. We conduct experiments using Meissonic, SD XL [29], FLUX.1-schnell [16], and SD-3.5-Large [40] on Challengebench, and present the results in Table 4. We observe that Meissonic’s improvement over SD XL on Challengebench (0.1838 $\rightarrow$ 0.2116) was greater than its improvement on HPD v2 (0.2888 $\rightarrow$ 0.2957) in Meissonic’s original paper, suggesting that MGT is more robust on challenging prompts. The reduced scores of FLUX.1-schnell and SD-3.5-Large on Challengebench indicate that it provides a robust measure of generalization. We achieve winning rates of approximately 70% using combination of design choices compared to vanilla sampling in Table 11. More experiments please see Appendix 9.9.
@@ 7.1 Benchmark
As detailed in Sec. 5, we initially generated 150,000 images using SD XL Base v1.0. We then filtered out the lowest-performing 1,500 prompts based on HPS v2 metrics. Finally, through a combination of manual filtering and GPT-4o filtering, we arrived at a final set of 220 semantically correct and complete prompts.
@@ 9.1 Additional Experiments of Different Noise Schedule
As a supplement to Sec. 3.1, we present the experimental results for the number of sampling steps $N=48$ and $N=32$, conducted on the HPD v2 Subset, in Figs 14 and 15, respectively. We further tested the performance of the noise schedule $1-t^{0.6}$ at $N=64$ on the full HPD v2 benchmark and presented the result in Table 7. Table 7 shows that $1-t^{0.6}$ outperforms $\mathrm{cos}(\frac{\pi t}{2})$ on the vast majority of metrics and subsets within the HPD v2, despite the potential risk of overfitting introduced by this exhaustive search approach on the hyperparameter $\rho$.
@@ 9.8 Additional Experiments on Challengebench
Due to space limitations in the main text, we present the experiments on Challengebench related to our proposed strategy in Table 15. From Table 15, it can be concluded that all of our strategies are effective.

### v2

@@ Abstract
Text-to-image diffusion models (DMs) develop at an unprecedented pace, supported by thorough theoretical exploration and empirical analysis. Unfortunately, the discrepancy between DMs and autoregressive models (ARMs) complicates the path toward achieving the goal of unified vision and language generation. Recently, the masked generative Transformer (MGT) serves as a promising intermediary between DM and ARM by predicting randomly masked image tokens (i.e., masked image modeling), combining the efficiency of DM with the discrete token nature of ARM. However, we find that the comprehensive analyses regarding the inference for MGT are virtually non-existent, and thus we aim to present positive design choices to fill this gap. We propose and redesign a set of enhanced inference techniques tailored for MGT, providing a detailed analysis of their performance. Additionally, we explore several DM-based approaches aimed at accelerating the sampling process on MGT. Extensive experiments and empirical analyses on the recent SOTA MGT, such as MaskGIT and Meissonic lead to concrete and effective design choices, and these design choices can be merged to achieve further performance gains. For instance, in terms of enhanced inference, we achieve winning rates of approximately 70% compared to vanilla sampling on HPS v2 with Meissonic-1024$\times$1024.
@@ 1 Introduction
Figure 2: The complete sampling pipeline of MaskGIT and Meissonic, and how our proposed effective and specific design choices integrate into that sampling pipeline. Specifically, TomeMGT and SCQ act on the Transformer to reduce inference latency and memory usage, respectively. Meanwhile, noise regularization and differential sampling enhance inference by correcting the probability distribution applied for sampling. Additionally, masked Z-Sampling is a rescheduling technique that significantly improves the quality of synthesized images through the forward-inversion operator (i.e., sampling and backtracking alternatively).
@@
To fill this gap, this paper focuses on the inference phase of MGT, with the primary objective of identifying design choices that enhance visual quality and the secondary goal of achieving efficient sampling through empirical analysis in high-resolution image synthesis scenarios. First, we elucidate how well-known training-free methods in DM can be applied to MGT, and the redesign required to ensure their effectiveness. As illustrated in Fig. 2 and Table 1, the sampling process of MGT bears significant similarity to that of DM, making it intuitive to adapt those algorithms from DM to MGT. We explore DPM-Solver [26, 25], TomeSD [4], and Z-Sampling [2] in this context, but find that all three algorithms require specific modifications to align with the characteristics of MGT in order to reduce NFE, accelerate inference, or achieve performance improvements. Given an example of Z-Sampling, we find that implementing DDIM Inversion [28], as used in DM, with random masking is ineffective unless masking is limited to low-confidence predicted tokens. In particular, our experimental outcomes indicate that among these three algorithms, DPM-Solver and TomeSD have relatively limited effects on MGT, whereas a rescheduling algorithm like Z-Sampling yields notable performance gains. Furthermore, we investigate the noise schedule in MGT, akin in EDM [16], highlighting that the cosine noise schedule is suboptimal under certain conditions. These findings suggest that inconsistencies between the training and inference mechanisms of DM and MGT lead to enhanced inference algorithms on DM not necessarily being effective for MGT.
@@
Second, we take a look at the probability distribution generated by the Transformer, leading to the development of several “cheap” (i.e., w/o significant computational overhead) yet effective distribution correction algorithms, including noise regularization and (low-entropy) differential sampling. To be specific, noise regularization dynamically applies (Gaussian) perturbations to the backbone output based on timesteps before applying softmax, aiming to enhance the diversity of synthesized images. Differential sampling, on the other hand, calculates the Kullback-Leibler (KL) divergence between the outputs of two adjacent time steps and resamples tokens with excessively similar Transformer outputs, thereby avoiding information redundancy and enhancing the visual quality.
@@
Noise Schedule (Section 3.1) $\sigma_{t}$ $\sqrt{e^{at^{2}+bt}-1}$ (VP-SDE [41]) $t$ (flow matching [24]) $(\sigma_{\textrm{max}}^{\frac{1}{\rho}}+t(\sigma_{\textrm{min}}^{\frac{1}{\rho}}-\sigma_{\textrm{max}}^{\frac{1}{\rho}}))^{\rho}$ (EDM [16]) N/A, and predicts one token per iteration $\mathrm{cos}\left(\frac{\pi t}{2}\right)$ $(1-t)^{\rho}$ or $1-t^{\rho}$
@@
Fourth, we evaluate the effectiveness of our proposed design choices on Meissonic using the HPD v2 benchmark, employing various metrics such as ImageReward [52], HPS v2 [50], PickScore [11, 21], and AES [19]. Similarly, we assess the performance of our strategy on MaskGIT using traditional metrics, including IS [35] and FID [11]. As shown in Fig. 1, these strategies lead to a significant improvement in the visual quality of synthesized images from Meissonic-512$\times$512 and Meissonic-1024$\times$1024.
@@ 3.3 Noise Regularization
where $\mathbf{v}_{i}$ is the output of $f_{\theta}$ and the yellow part stands for noise regularization. Noise regularization introduces randomness (see Appendix 9.2 for more details) into the sampling process, enhancing the diversity of predicted tokens.
@@
In our definition, the hyperparameter $\mathbf{I}_{t}$ is a function $\mathbf{I}(t)$ that represent the standard deviation of the noise, which varies across different timesteps $t$. To determine the empirically optimal $\mathbf{I}(t)$, we test various curves and calculate the winnowing rates of four metrics—PickScore, HPS v2, AES, and ImageReward—relative to vanilla sampling. The results are presented in Fig. 7. We find that $\mathbf{I}(t)$ works best when set to $|\cos(\pi t)|$, which reaches larger values as $t$ approaches 00 or $1$ and lower values in the middle. This observation is interesting, as injecting noise during the end-of-sampling phase significantly degrades inference performance on DM, while having positive impact on MGT. The results on HPD v2 can be found in Appendix 9.4.
@@ 3.4 Differential Sampling
T2I-Compbench Attribute Binding Object Relationship
(Meissonic) Color (↑) Texture (↑) 2D-Spatial (↑) 3D-Spatial (↑) Non-Spatial (↑) Complex (↑)
Vanilla Sampling 0.5540 0.4858 0.1809 0.3381 0.3037 0.2942
Noise Regularization 0.5682 0.4937 0.1932 0.3687 0.3053 0.2976
Differential Sampling 0.5458 0.4561 0.1515 0.3598 0.3041 0.2994
Masked Z-Sampling 0.5451 0.5106 0.1738 0.3642 0.3033 0.2913
Table 4: Comparison of the combination of our methods on T2I-Compbench [14] with Meissonic-1024$\times$1024.
GenEval Single Object (↑) Two Object (↑) Counting (↑) Colors (↑) Position (↑) Color Attr (↑) Avg. (↑)
Vanilla Sampling 91.25% 54.61% 37.50% 79.91% 9.00% 15.00% 47.87%
Noise Regularization 95.00% 51.52% 41.25% 80.85% 5.00% 20.00% 48.93%
Differential Sampling 92.50% 54.44% 33.75% 79.66% 13.00% 15.00% 48.05%
Masked-Z Sampling 93.75% 55.56% 36.25% 84.04% 9.00% 26.00% 50.77%
Table 5: Comparison of the combination of our methods on GenEval [9] with Meissonic-1024$\times$1024.
@@ 4.2 Introducing TomeSD into MGT
Token merging is designed to enhance efficiency by first merging tokens after the linear layers and subsequently unmerging them. Applying token merging to accelerate inference is natural, as the backbone of MGT is a Transformer. Unfortunately, Meissonic has only 1024 tokens, fewer than 4096 tokens in the attention layer of SD XL. As is well known, the computational complexity of attention increases exponentially with the number of tokens, and a smaller number of tokens reduces the potential benefits of token merging, resulting in an insignificant effect of our implemented TomeMGT as observed in our experiments. Therefore, we focus on the challenges of applying TomeSD, which has demonstrated effective on SD XL, to MGT to achieve accelerated inference, along with the corresponding application scenarios.
@@ 4.3 Momentum-based Solver
5 Additional Evaluations
MaskGIT-512$\times$512 IS (↑) FID (↓) Prec. (↑) Recall (↑) sFID (↓)
Vanilla Sampling 291.30 14.55 0.8403 0.165 39.26
Noise Regularization 294.72 14.18 0.8666 0.184 36.90
Differential Sampling 291.30 14.35 0.8658 0.166 39.27
Masked-Z Sampling 298.12 12.87 0.8842 0.199 33.58
Table 7: Comparison results of MaskGIT on traditional metrics.
To substantiate the generalization ability of the enhanced inference algorithm proposed by us, we conduct experiments on more MGTs (i.e., Meissonic-512$\times$512 and MaskGIT-512$\times$512) as well as more benchmarks (i.e., HPD v2, GenEval and T2I-Compbench). The experimental results on Meissonic-1024$\times$1024 within T2I-Compbench [14], GenEval [9], and HPD v2 are summarized in Table 4, Table 5, and Fig. 10, respectively. These results clearly highlight the effectiveness of all proposed design strategies. Specifically, when integrated with the other two strategies, masked Z-Sampling achieves winning rates of approximately 70% compared to vanilla sampling on HPD v2. For clarity and due to space limitations in the main paper, the remain experimental results for Meissonic-512$\times$512 and Meissonic-1024$\times$1024 are provided in Appendix 9.9. We further validate the significant performance of our proposed design strategies on class-to-image (C2I) MaskGIT [5]. As presented in Table 7, all strategies contribute to performance gains, with masked Z-Sampling yielding the most notable improvements. Moreover, we demonstrate that these strategies are capable of refining token distributions even on ARM (e.g., LlamaGen [44]). For further details, please refer to Appendix 9.11.
@@ 7.1 Benchmark
This benchmark was proposed by us, with the detailed prompt collection methodology described in Appendix 9.10. The core objective of this benchmark is to select challenging prompts, thereby enabling an exploration of the performance limits of generative models. We initially generated 150,000 images using SD XL Base v1.0. We then filtered out the lowest-performing 1,500 prompts based on HPS v2 metrics. Finally, through a combination of manual filtering and GPT-4o filtering, we arrived at a final set of 220 semantically correct and complete prompts.
GenEval.
GenEval is an object-focused framework designed to evaluate compositional image properties, including object co-occurrence, position, count, and color. This benchmark utilizes state-of-the-art object detection models to assess text-to-image generation tasks, ensuring strong alignment with human agreement. Additionally, other discriminative vision models can be integrated into this pipeline to further verify attributes such as object color. Notably, this benchmark comprises 550 prompts, each of which is straightforward and easy to understand.
T2I-Compbench.
T2I-Compbench is a benchmark similar to GenEval, designed as a comprehensive evaluation framework for open-world compositional text-to-image synthesis. It comprises 6,000 compositional text prompts, categorized into three main groups (attribute binding, object relationships, and complex compositions) and further subdivided into six subcategories: color binding, shape binding, texture binding, spatial relationships, non-spatial relationships, and complex compositions.
@@ 9.1 Additional Experiments of Different Noise Schedule
As a supplement to Sec. 3.1, we present the experimental results for the number of sampling steps $N=48$ and $N=32$, conducted on the HPD v2 Subset, in Figs 13 and 14, respectively. We further tested the performance of the noise schedule $1-t^{0.6}$ at $N=64$ on the full HPD v2 benchmark and presented the result in Table 9. Table 9 shows that $1-t^{0.6}$ outperforms $\mathrm{cos}(\frac{\pi t}{2})$ on the vast majority of metrics and subsets within the HPD v2, despite the potential risk of overfitting introduced by this exhaustive search approach on the hyperparameter $\rho$.
9.2 Analysis of Noise Regularization and Differential Sampling
Here, we present empirical arguments for why noise regularization works. As shown in Fig. 15 (Left), vanilla sampling often results in horizontal bars of the same color as $t$ approaches 0 and vertical bars of the same color as $t$ approaches 1, indicating the redundancy of its sampling process. To be specific, we calculate the mean values of the absolute differences between adjacent horizontal bars and vertical bars: standard sampling is $\frac{43.2}{255}$ and $\frac{27.0}{255}$, while noise regularization is $\frac{45.6}{255}$ and $\frac{35.3}{255}$. The increased variability in bar intensities under noise regularization—evidenced by higher mean differences—quantitatively confirms the method’s efficacy in reducing same-color bar artifacts. Interestily, as presented in Fig. 15 (Right), noise regularization modifies the entropy of the distribution during the sampling phase, making it more inclined toward a “U”-shaped structure, which is similar to the shape of $|\cos(\pi t)|$.
@@ 9.9 Verification of Combinations of Algorithms
9.10 ChallengeBench
We further analyze the performance of MGT on challenging prompts. We synthesize 150k images using SD XL [31], computed HPS v2 scores, and selected 1.5k prompts with the lowest scores. After manual and GPT-4o [1] selection, we form Challengebench with 220 semantically sound prompts. We conduct experiments using Meissonic, SD XL [31], FLUX.1-schnell [18], and SD-3.5-Large [42] on Challengebench, and present the results in Table 17. We observe that Meissonic’s improvement over SD XL on Challengebench (0.1838 $\rightarrow$ 0.2116) was greater than its improvement on HPD v2 (0.2888 $\rightarrow$ 0.2957) in Meissonic’s original paper, suggesting that MGT is more robust on challenging prompts. Furthermore, our inference-enhancing algorithms including noise regularization, masked Z-Sampling, and differential sampling continue to have significant performance gains against Meissonic on this benchmark.
9.11 Additional Experiments on LlamaGen
Our findings reveal that differential sampling and noise regularization are independent of the inversion operator (masked Z-Sampling cannot implement since lack the inversion operator) and can be directly implemented on ARM. We conducted experiments using LlamaGen [44], with the corresponding results shown in Table 18 and Table 19. Our experiments reveal that these two strategies can significantly improve LlamaGen’s generative performance while incurring almost no additional computational overhead. This finding highlights a high alignment in predictive objectives between MGT and ARM.
GenEval Single Object (↑) Two Object (↑) Counting (↑) Colors (↑) Position (↑) Color Attr (↑) Avg. (↑)
(LlamaGen-512$\times$512)
Vanilla Sampling 20.00% 9.09% 1.25% 15.96% 11.00% 0.00% 9.55%
Noise Regularization 20.00% 9.09% 2.50% 13.83% 16.00% 1.00% 10.40%
Differential Sampling 20.00% 12.12% 2.50% 14.26% 14.00% 2.00% 10.81%
Table 18: Comparison of our proposed methods on GenEval [9] with LlamaGen-512$\times$512.
LlamaGen-256$\times$256 IS (↑) FID (↓) Prec. (↑) Recall (↑) sFID (↓)
Vanilla Sampling 316.68 13.22 0.8723 0.215 18.10
Noise Regularization 317.41 13.02 0.8741 0.207 17.44
Differential Sampling 317.80 12.87 0.8732 0.206 15.87
Table 19: Comparison of our proposed methods on the traditional metrics with LlamaGen-256$\times$256.